if a 10 ft tall statue casts a 20 ft long shadow, then how tall is a tent that casts a 10 ft shadow?

Answer by jim_thompson5910(35256)

(height of statue)/(length of statue's shadow) = (height of tent)/(length of tent's shadow)
(10)/(20) = (x)/(10)
10/20 = x/10
1/2 = x/10
1*10 = 2*x
10 = 2x
2x = 10
x = 10/2
x = 5
The tent is 5 feet tall.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------